OpenStudy (anonymous):

Solve the system algebraically: 32x^2 + 9y^2 = 324 3y^2 - x^2 = 3

OpenStudy (displayerror):

You can solve by substitution. Either solve for x in terms of y or y in terms of x using one equation, then substitute that into the other equation to solve. Using the second equation, if we solve for x first: \[3y^2-x^2=3\] \[x^2 = 3y^2-3\] Now we have x^2 in terms of y^2. Notice that there is also an x^2 term in the first equation. We can then substitute the expression for x^2 into the first equation wherever we see x^2: The first equation: \[32x^2+9y^2=324\] Substituting in for x^2: \[32\left( 3y^2-3 \right) + 9y^2 = 324\] Distribute the 32: \[96y^2 - 96 + 9y^2 = 324\] Combine like terms to solve for y. Once we find y, we can plug that into our expression for x^2 to solve for x \[x^2 = 3y^2 -3\]

OpenStudy (anonymous):

@DisplayError I got \[y = \pm4 \] and \[y = \pm6.7\] Is that what I was supposed to get or did you get something different?

OpenStudy (anonymous):

That second equation is supposed to be \[x = \pm6.7\]

OpenStudy (displayerror):

You're close! Remember, we need to solve for y, not y^2. From this equation: \[96y^2 - 96 + 9y^2 = 324\] We can combine the y^2 terms and move the 96 to the other side: \[105y^2 = 420\] Now solve for y, not y^2.

OpenStudy (anonymous):

Oh it would be y = +- 2!

OpenStudy (displayerror):

Yep! Now plug into the equation for x^2 to solve for x: \[x^2 = 3y^2 -3\]

OpenStudy (anonymous):

And x would be \[x=\pm3\] so does that mean it intersects at 4 points?

OpenStudy (displayerror):

Yep! You should get the points (2, 3), (2, -3), (-2, 3), and (-2, -3), so four points of intersection.

OpenStudy (displayerror):

I've attached a plot of the two polynomials to help you see how it looks like